# MaDiS: Taming Masked Diffusion Language Models for Sign Language Generation

### v1

@@ Abstract
Sign language generation (SLG) aims to translate written texts into expressive sign motions, bridging communication barriers for the Deaf and Hard-of-Hearing communities. Recent studies formulate SLG within the language modeling framework using autoregressive language models, which suffer from unidirectional context modeling and slow token-by-token inference. To address these limitations, we present MaDiS, a masked-diffusion-based language model for SLG that captures bidirectional dependencies and supports efficient parallel multi-token generation. We further introduce a tri-level cross-modal pretraining scheme that jointly learns from token-, latent-, and 3D physical-space objectives, leading to richer and more grounded sign representations. To accelerate model convergence in the fine-tuning stage, we design a novel unmasking strategy with temporal checkpoints, reducing the combinatorial complexity of unmasking orders by over $10^{41}$ times. In addition, a mixture-of-parts embedding layer is developed to effectively fuse information stored in different part-wise sign tokens through learnable gates and well-optimized codebooks. Extensive experiments on CSL-Daily, Phoenix-2014T, and How2Sign demonstrate that MaDiS achieves superior performance across multiple metrics, including DTW error and two newly introduced metrics, SiBLEU and SiCLIP, while reducing inference latency by nearly 30%. Code and models will be released on our project page.
Figure 1: We propose MaDiS, a novel sign language generation approach built upon masked diffusion language models (MDLMs) [45]. (a) A sign tokenizer discretizes continuous sign motions into part-wise tokens [96]. (b) Conventional autoregressive language models generate tokens in a left-to-right manner, limiting utilization of contexts and inference efficiency. (c) The emerging MDLMs model token distributions with bidirectional contexts and enable parallel multi-token sampling during inference. (d) MaDiS achieves SOTA performance across multiple benchmarks [18, 94, 11] while reducing inference latency by nearly 30%.
@@ 1 Introduction
Sign language is the primary communication mode for Deaf and Hard-of-Hearing (DHH) individuals, yet in predominantly oral societies they are often encouraged to rely on text instead of their preferred sign language [85]. Research in sign language translation (SLT, sign-to-text) and generation (SLG, text-to-sign) aims to improve the accessibility of sign language, thereby mitigating the communication gap between the hearing and DHH communities. While SLT has seen significant progress [37, 23, 14, 31, 76, 28, 21, 98, 46], SLG remains relatively underexplored.
@@
Typical SLG works [64, 5, 19, 67, 73] formulate the task as a visual content generation problem, using either GANs [61] or diffusion models [56]. Recently, motivated by the linguistic character of sign language and the generalizability of large language models (LLMs), a new research direction has emerged that models SLG through tokenization (Figure 1a) and autoregressive language models (ARLMs) [96, 84, 87, 20, 17] (Figure 1b). However, ARLMs have two main limitations: their left-to-right generation order restricts the use of bidirectional contexts and their one-token-at-a-time decoding introduces an inference bottleneck.
@@
To address such problems, we propose MaDiS (Figure 1c), a novel SLG approach built upon an emerging and promising language modeling paradigm, the masked diffusion language model (MDLM) [45, 82, 86, 44]. Unlike conventional ARLMs, MDLMs are based on masked diffusion models [62, 58, 47], which introduce a forward data masking process and train a mask predictor to approximate the reverse denoising process. This mechanism enables MDLMs to construct model distributions with bidirectional dependencies, allowing them to capture richer contextual information. During the reverse process, multiple tokens can be sampled simultaneously, thereby improving inference efficiency.
@@
Pretraining plays a crucial role in sign language understanding models [37, 29, 95, 77, 92, 34], yet pretraining for SLG models remains largely unexplored. MDLMs are typically pretrained in the token space using a mask-then-predict framework [45], where a subset of tokens is randomly masked and subsequently predicted by the model. Although such token-space objective is well-established in natural language processing, recent advances in self-supervised learning [36, 6, 4] have demonstrated that latent-space pretraining is often more effective in a wide range of vision tasks, including image and video understanding [43, 6] and world modeling [93, 3]. Moreover, sign language, as a visual language, conveys most of its semantics through body and hand motions in the 3D physical space. Motivated by the advantages of different representation spaces, we propose a tri-level cross-modal pretraining method in which the model learns to predict masked text and sign tokens, latent features, and 3D sign motions simultaneously. This challenging and comprehensive pretraining objective encourages the model to learn rich and grounded sign representations and yields significant performance improvements in the downstream SLG task.
@@
After pretraining, we perform supervised fine-tuning conditioned on unmasked text tokens. The original MDLM can be viewed as an any-order ARLM [65]. Due to its unconstrained confidence-based unmasking strategy, the number of possible unmasking orders is enormous. For example, generating 100 tokens in 25 steps results in approximately $10^{123}$ possible orders. Although this property enables MDLMs to perform well in data-constrained scenarios, it also leads to significantly slower convergence, often requiring more training epochs [52]. To address this issue, we propose a novel unmasking strategy with temporal checkpoints. Inspired by the effectiveness of VAR [70], which generates images progressively from low to high resolution, we insert checkpoints at noise levels 0.75 and 0.5, corresponding to temporal resolutions of $T/4$ and $T/2$. This design dramatically prunes unmasking orders by over $10^{41}$ times, thereby accelerating convergence during training. Furthermore, to better integrate part-wise sign tokens [96], we propose a plug-and-play mixture-of-parts (MoP) embedding layer. Unlike the previous SOTA method [96] that simply averages token embeddings from different body parts, our MoP embedding layer is built upon well-optimized VQ-VAE codebooks, and part-wise embeddings are dynamically fused through learnable gates. In summary, our contributions are as follows:
@@
  • We propose MaDiS, a novel SLG approach built upon MDLMs that enable bidirectional context modeling and parallel multi-token generation.
@@
  • A tri-level cross-modal pretraining framework is developed that jointly learns from token, latent, and 3D physical spaces, enabling the model to learn richer and more grounded sign representations.
@@
  • To accelerate training convergence, we design a novel token unmasking strategy with temporal checkpoints, which significantly reduces unmasking order complexity. To enhance the integration of part-wise sign tokens, we introduce an MoP embedding layer that is conditioned on VQ-VAE codebooks and controlled by learnable gates.
@@
  • Extensive experiments on widely adopted benchmarks [18, 94, 11] demonstrate that MaDiS achieves SOTA performance across multiple metrics, including DTW error and two newly introduced metrics, SiBLEU and SiCLIP, while reducing inference latency by $\sim$30% (Figure 1d).
@@
2 Related Work
Sign Language Generation (SLG). SLG aims to translate texts into sign language outputs represented in various forms, such as skeletons [60, 67, 1, 79, 68], RGB videos [73, 61, 53], and 3D motions [96, 97, 84, 54]. Although real-person videos are more realistic, developing end-to-end sign video generation models is challenging due to the high dimensionality of RGB video data. Therefore, we focus on generating sign motions, which can serve as an intermediate representation for subsequent sign video synthesis [64, 63]. Motivated by the linguistic nature of sign language, recent studies [87, 17, 96, 84] formulate SLG within a language modeling framework. However, these approaches rely on ARLMs, which are inherently limited by their left-to-right generation process. In this work, we explore emerging MDLMs for SLG, which model token distributions bidirectionally and enable efficient parallel generation.
Masked Diffusion Language Model (MDLM). Current LLMs are mostly built upon ARLMs. Although ARLMs demonstrate strong capability, their left-to-right generation order limits both efficiency and context modeling. Recently, MDLMs [45, 44, 24] have emerged as a new language modeling paradigm, showing promising capability and scalability, even surpassing ARLMs on the reversal poem completion task. MDLMs have also been successfully applied to diverse domains, including multimodal understanding [82], protein design [74], and robotics [75]. A related line of work in human motion generation [27, 26, 51] adopts a similar masked generation approach [13], but their training objectives are determined heuristically and lack the rigorous theoretical foundation of MDLMs. In this work, we present the first effort to introduce MDLMs into SLG and propose a temporal-checkpoint unmasking strategy to improve training efficiency of MDLMs [52].
Pretraining for Sign Language Models. Pretraining plays a crucial role in sign language understanding tasks such as recognition and translation. Existing works can be categorized into alignment-based methods [34, 92, 35, 83, 15], which pretrain by aligning signs and texts; reconstruction-based methods [29, 77, 25, 57, 91, 90], which pretrain by reconstructing sign inputs; hybrid methods [95]; and next-token prediction methods [37, 88]. However, an important oversight is that all these works focus solely on understanding tasks. Our work represents the first attempt to pretrain sign language generation models through a cross-modal and multi-level strategy that jointly predicts sign and text tokens, latent features, and 3D sign motions, thereby effectively injecting sign-aware knowledge into MDLMs.
@@ 3 Methodology
An overview of the proposed MaDiS is illustrated in Figure 2. In the first stage, both text and sign tokens are randomly masked according to a masking ratio $t$, and the MDLM is pretrained with objectives defined across multiple representation spaces (token, latent, and physical) to exploit the benefits from multi-level sign representations. In the subsequent fine-tuning stage, the MDLM keeps text tokens unmasked and learns to generate sign tokens using the proposed unmasking with temporal checkpoints (UTC) strategy and the mixture-of-parts (MoP) embedding layer.
@@ 3.1 Preliminaries
Sign Tokenizer. We adopt the open-sourced SOTA decoupled tokenizer [96] to convert continuous sign motions into discrete tokens. It consists of three parallel VQ-VAEs [72] responsible for different body parts. Each VQ-VAE includes an encoder $\mathcal{E}_{p}$, a decoder $\mathcal{D}_{p}$, and a codebook $\mathcal{C}_{p}=\{{\mathbf{c}_{p}^{i}}\}_{i=1}^{N_{c}}$, where $p\in\{b,l,r\}$ represents the upper body, left hand, and right hand, respectively, and $N_{c}$ is the codebook size. Using this tokenizer, a $T$-frame sign motion sequence $\mathbf{S}\in\mathbb{R}^{T\times d_{s}}$, where $d_{s}=133$ denotes the number of SMPL-X parameters [49], is decomposed as three part-wise token sequences $x_{p}=\{x_{p}^{i}\}_{i=1}^{L_{s}}$, where $x_{p}^{i}\in\{1,\dots,N_{c}\}$ and $L_{s}$ denotes the token sequence length.
Masked Diffusion Language Model. MDLMs [45, 44, 82] introduce a new language modeling paradigm based on a forward-reverse discrete diffusion process. In the forward process, the original token sequence $x_{0}$ is corrupted with mask tokens $|\mathrm{M}|$ according to the time (noise level) $t\in[0,1]$. The conditional distribution of the masked sequence $x_{t}$ is defined as $q_{t|0}(x_{t}|x_{0})=\prod_{i=1}^{L}q_{t|0}(x_{t}^{i}|x_{0}^{i})$, where
@@
A rigorous proof in [47] shows that $q_{0|t}(x_{s}^{i}|x_{t})=p(x_{0}^{i}|x_{t}^{\mathrm{UM}})$, where $x_{t}^{\mathrm{UM}}$ denotes the unmasked tokens in $x_{t}$ that are identical to those in $x_{0}$. This implies that the time condition $t$ can be omitted from the model inputs, leading to a simplified implementation of MDLM in which the model only needs to predict masked tokens conditioned on the unmasked ones.
@@ 3.2 Tri-Level Cross-Modal Pretraining
Existing works [12, 71] reveal that the knowledge of LMs mainly comes from pretraining. As shown in Figure 2, we pretrain the MDLM at three spaces, enabling it to absorb knowledge from sign representations of different levels.
Token Space. At the first level, we pretrain the MDLM by predicting masked tokens. Following [45], at each training iteration, we uniformly sample a noise level (mask ratio) $t\sim U(0,1)$. Denoting text tokens as $x_{e}=\{x_{e}^{i}\}_{i=1}^{L_{e}}$ and an end-of-sentence token as $|\mathrm{eos}|$, the input sequence is formed by concatenating text tokens, $|\mathrm{eos}|$, and sign tokens: $x_{0}=\mathrm{cat}(x_{e},|\mathrm{eos}|,x_{\{p\}},|\mathrm{eos}|)=\{x_{0}^{i}\}_{i=1}^{L_{e}+L_{s}+2}$. For simplicity, we use $\{p\}$ to denote all part-wise sign tokens, since they are masked and unmasked simultaneously once their index is selected. Each index $i$ is selected with a probability of $t$, and its corresponding token is masked to obtain $x_{t}$. The token-space loss is defined as a cross-entropy loss [45]:
$\mathcal{L}_{tok}=-\mathbb{E}_{t,x_{0},x_{t}}\left[\frac{1}{t}\sum_{i\in I_{m}}\log p_{\theta}(x_{0}^{i}\mid x_{t})\right],$ (3)
@@
Latent Space. Inspired by the effectiveness of predicting embeddings (features), e.g., joint-embedding predictive architectures (JEPAs) [36, 2], in numerous vision tasks [6, 3, 93], we extend this idea to MDLM pretraining. In contrast to the original JEPAs that train an additional target encoder to extract features, we directly leverage the well-trained codebooks $\mathcal{C}_{p}$, whose code embeddings are natural sign representations, as the targets. More specifically, let the last hidden states of the MDLM corresponding to masked sign tokens be $\mathbf{H}\in\mathbb{R}^{|I_{s}|\times d}$, where $I_{s}$ denotes the index set of masked sign tokens. We use three MLPs to map $\mathbf{H}$ into the spaces of the codebooks: $\mathbf{H}_{p}=\mathrm{MLP}_{p}(\mathbf{H})\in\mathbb{R}^{|I_{s}|\times d_{c}}$, where $d_{c}$ denotes the dimension of each code embedding. The target embeddings are then collected from the corresponding code embeddings: $\mathbf{\hat{H}}_{p}=\mathrm{cat}(\{\mathbf{c}^{n_{i}}_{p}\mid i\in I_{s},1\leq n_{i}\leq N_{c}\})$, where $n_{i}$ denotes the code index of the $i$-th token. Finally, the MDLM learns to predict the code embeddings with a smoothed L1 loss:
$\mathcal{L}_{lat}=\sum_{p}\mathrm{Smooth\_L1}(\mathbf{H}_{p},\mathbf{\hat{H}}_{p}).$ (4)
Physical Space. Inspired by the visual–language nature of sign languages, we incorporate 3D sign motion reconstruction as an additional pretraining objective in the physical space. More specifically, we reuse the well-trained VAE decoder $\mathcal{D}_{p}$, which is frozen and used to reconstruct sign motions from the mapped hidden states $\mathbf{H}_{p}$. Denoting the ground-truth part-wise sign motions at the masked indices as $\mathbf{\hat{S}}_{p}$, the objective function in the physical space is defined as:
$\mathcal{L}_{phy}=\sum_{p}\mathrm{Smooth\_L1}(\mathcal{D}_{p}(\mathbf{H}_{p}),\mathbf{\hat{S}}_{p}).$ (5)
Unlike existing works [29, 95, 57] that mask and reconstruct sign poses, we mask tokens but reconstruct poses (motions), which is more challenging and encourages the model to learn more grounded sign representations. The overall pretraining loss is defined as $\mathcal{L}_{pre}=\mathcal{L}_{tok}+\mathcal{L}_{lat}+\mathcal{L}_{phy}$.

### v2

@@ Abstract
Sign language generation (SLG) aims to translate written texts into expressive sign motions, bridging communication barriers for the Deaf and Hard-of-Hearing communities. Recent studies formulate SLG within the language modeling framework using autoregressive language models, which suffer from unidirectional context modeling and slow token-by-token inference. To address these limitations, we present MaDiS, a masked-diffusion-based language model for SLG that captures bidirectional dependencies and supports efficient parallel multi-token generation. We further introduce a tri-level cross-modal pretraining scheme that jointly learns from token-, latent-, and 3D physical-space objectives to leverage complementary, multi-level sign representations. To accelerate model convergence in the fine-tuning stage, we design a novel unmasking strategy with temporal checkpoints, which restructures generation in a coarse-to-fine manner and reduces the combinatorial complexity of unmasking orders by over $10^{41}$ times. In addition, a mixture-of-parts embedding layer is developed to effectively fuse information stored in different part-wise sign tokens through a learnable gate and well-optimized codebooks. Extensive experiments on CSL-Daily, Phoenix-2014T, and How2Sign demonstrate that MaDiS achieves superior performance across multiple metrics, including DTW error and two newly introduced metrics, SiBLEU and SiCLIP, while delivering a 40% higher throughput. Code and models will be publicly released.
Figure 1: We propose MaDiS, a novel sign language generation approach built upon masked diffusion language models (MDLMs) [llada]. (a) A sign tokenizer discretizes continuous sign motions into part-wise tokens [zuo2025soke]. (b) Conventional autoregressive language models generate tokens in a left-to-right manner, limiting utilization of contexts and inference efficiency. (c) The emerging MDLMs model token distributions with bidirectional contexts and enable parallel multi-token sampling during inference. (d) MaDiS achieves SOTA performance across multiple benchmarks [duarte2021how2sign, csl-daily, 2014T] while delivering a 40% higher throughput.
@@ 1 Introduction
Sign language is the primary communication mode for Deaf and Hard-of-Hearing (DHH) individuals, yet in predominantly oral societies they are often encouraged to rely on text instead of their preferred sign language [yin2021including]. Research in sign language translation (SLT, sign-to-text) and generation (SLG, text-to-sign) aims to improve the accessibility of sign language, thereby mitigating the communication gap between the hearing and DHH communities. While SLT has seen significant progress [li2025uni-sign, gong2024llms, chentwo, jang2025lost, wongsign2gpt, guo2025bridging, fish2025geo-sign, gan2025mixsigngraph, yao2023sign], SLG remains relatively under-explored.
Early SLG approaches typically rely on glosses, the written form of signs, as an intermediate representation and formulate SLG as a text-to-gloss-to-sign pipeline [zuo2024simple, saunders2022signing, saunders2020progressive, stoll2020text2sign]. While glosses encode rich linguistic information, the performance of gloss-based methods is constrained by the quality of the text-to-gloss module: due to the high cost of gloss annotations, gloss-based datasets are often limited in scale, leading to degraded text-to-gloss performance. For example, Gemini-generated glosses exhibit an error rate of over 60% [guo2025bridging]. In addition, such cascaded pipelines incur higher inference latency.
@@
Given these limitations, recent efforts have increasingly adopted gloss-free SLG, directly generating signs from text. Typical works [Shi_2025_ICCV, baltatzis2024nsa, fang2023signdiff, wang2025signvip] formulate the task as a visual content generation problem, using either GANs [saunders2022signing] or diffusion models [latentdiff]. Recently, motivated by the linguistic character of sign language and the generalizability of large language models (LLMs), a new research direction has emerged that models SLG through tokenization (Fig.˜1a) and autoregressive language models (ARLMs) [zuo2025soke, yin-etal-2024-t2s, yu2025signavatars, fang2025signllm, fg2024signavatar] (Fig.˜1b). However, ARLMs have two main limitations: their left-to-right generation order restricts the use of bidirectional contexts and their one-token-at-a-time decoding introduces an inference bottleneck.
@@
To address such problems, we propose MaDiS (Fig.˜1c), a novel SLG approach built upon an emerging and promising language modeling paradigm, the masked diffusion language model (MDLM) [llada, yang2025mmada, llada-v, smdm]. Unlike conventional ARLMs, MDLMs are based on masked diffusion models [shi2024simplified, sahoo2024simple, ou2025your], which introduce a forward data masking process and train a mask predictor to approximate the reverse denoising process. This mechanism enables MDLMs to construct model distributions with bidirectional dependencies, allowing them to capture richer contextual information. During the reverse process, multiple tokens can be sampled simultaneously, thereby improving inference efficiency.
@@
Pretraining plays a crucial role in sign language understanding models [li2025uni-sign, hu2023signbert+, zhou2025scaling, Wong_2025_signrep, zhou2023gloss, jiao2024visual], yet pretraining for SLG models remains largely unexplored. MDLMs are typically pretrained in the token space using a mask-then-predict framework [llada], where a subset of tokens is randomly masked and subsequently predicted by the model. Although such token-space objective is well-established in natural language processing, recent advances in self-supervised learning [jepa, v-jepa, baldassarre2025back] have demonstrated that latent-space pretraining is often more effective in a wide range of vision tasks, including image and video understanding [mo2024connecting, v-jepa] and world modeling [dino-wm, v-jepa2]. Moreover, sign language, as a visual language, conveys most of its semantics through body and hand motions in the 3D physical space. Motivated by the advantages of different representation spaces, we propose a tri-level cross-modal pretraining method in which the model learns to predict masked text and sign tokens, latent features, and 3D sign motions simultaneously. This challenging and comprehensive pretraining objective encourages the model to fully exploit complementary sign representations and yields significant performance improvements in the downstream SLG task.
@@
After pretraining, we perform supervised fine-tuning conditioned on unmasked text tokens. The original MDLM can be viewed as an any-order ARLM [seed-diffusion]. Due to its unconstrained confidence-based unmasking strategy, the number of possible unmasking orders is enormous. For example, generating 100 tokens in 25 steps results in approximately $10^{123}$ possible orders. Although this property enables MDLMs to perform well in data-constrained scenarios, it also leads to significantly slower convergence, often requiring more training epochs [prabhudesai2025diffusion]. To address this issue, we propose a novel unmasking strategy with temporal checkpoints. Inspired by the effectiveness of VAR [tian2024var], which generates images progressively from low to high resolution, we insert checkpoints composed of uniformly distributed tokens at noise levels 0.75 and 0.5, corresponding to temporal resolutions of $T/4$ and $T/2$. This design decomposes generation in a coarse-to-fine manner and dramatically prunes unmasking orders by over $10^{41}$ times, thereby accelerating convergence during training. Furthermore, to better integrate part-wise sign tokens [zuo2025soke], we propose a plug-and-play mixture-of-parts (MoP) embedding layer. Unlike the previous SOTA method [zuo2025soke] that simply averages token embeddings from different body parts, our MoP embedding layer is built upon well-optimized VQ-VAE codebooks, and part-wise embeddings are dynamically fused through a learnable gate. In summary, our contributions are as follows:
@@
  • We propose MaDiS, a novel gloss-free SLG approach built upon MDLMs that enable bidirectional context modeling and parallel multi-token generation.
@@
  • A tri-level cross-modal pretraining framework is developed that jointly learns from token, latent, and 3D physical spaces, enabling the model to exploit the benefits from complementary, multi-level sign representations.
@@
  • To accelerate training convergence, we design a novel token unmasking strategy with temporal checkpoints, which significantly reduces unmasking order complexity through a coarse-to-fine decomposition. To enhance the integration of part-wise sign tokens, we introduce an MoP embedding layer that is conditioned on VQ-VAE codebooks and controlled by a learnable gate.
@@
  • Extensive experiments on widely adopted benchmarks [duarte2021how2sign, csl-daily, 2014T] demonstrate that MaDiS achieves SOTA performance across multiple metrics, including DTW error and two newly introduced metrics, SiBLEU and SiCLIP, while enhancing inference throughput by 40% (Fig.˜1d).
2 Related Work
Sign Language Generation (SLG). SLG aims to translate texts into sign language outputs represented in various forms, such as skeletons [saunders2021mixed, tang2025discrete, arkushin2023ham2pose, xie2024g2p, tang2022gloss], RGB videos [wang2025signvip, saunders2022signing, signgen, fang2025stable, fang2023signdiff], and 3D motions [zuo2025soke, zuo2024simple, yin-etal-2024-t2s, yu2025signavatars, baltatzis2024nsa, bensabath2025text]. Although real-person videos are more realistic, developing end-to-end sign video generation models is challenging due to the high dimensionality of RGB video data. Therefore, we focus on generating sign motions, which can serve as an intermediate representation for subsequent sign video synthesis [Shi_2025_ICCV, shi2024pose, wang2025signaligner]. Motivated by the linguistic nature of sign language, recent studies [yu2025signavatars, fg2024signavatar, zuo2025soke, yin-etal-2024-t2s] formulate SLG within a language modeling framework. However, these approaches rely on ARLMs, which are inherently limited by their left-to-right generation process. In this work, we explore emerging MDLMs for SLG, which model token distributions bidirectionally and enable efficient parallel generation.
Masked Diffusion Language Model (MDLM). Current LLMs are mostly built upon ARLMs. Although they demonstrate strong capability, their left-to-right generation order limits both efficiency and context modeling. Recently, MDLMs [llada, bie2025llada2, smdm, gongscaling] have emerged as a new language modeling paradigm, showing promising capability, scalability, and efficiency. MDLMs have also been successfully applied to diverse domains, including multimodal understanding [yang2025mmada], protein design [wang2024diffusion], and robotics [wen2025llada-vla]. A related line of work in human motion generation [guo2024momask, guo2025snapmogen, pinyoanuntapong2025maskcontrol, xia2026signmask] adopts a similar masked generation approach [chang2022maskgit], but their training objectives are determined heuristically and lack the rigorous theoretical foundation of MDLMs. In this work, we present the first effort to introduce MDLMs into SLG and propose a temporal-checkpoint unmasking strategy to improve training efficiency.
@@ 2 Related Work
Pretraining for Sign Language Models. Pretraining plays a crucial role in sign language understanding tasks. Existing works can be categorized into alignment-based methods [jiao2024visual, zhou2023gloss, kim2025leveraging, ye2024improving, chen2025c2rl], which pretrain by aligning signs and texts; reconstruction-based methods [hu2023signbert+, Wong_2025_signrep, gueuwou2025shubert, rust-etal-2024-privacy, best, zhao2024masa], which pretrain by reconstructing sign inputs; hybrid methods [zhou2025scaling]; and next-token prediction methods [li2025uni-sign, zhang2024scaling]. However, an important oversight is that all these works focus solely on understanding tasks. Our work represents the first attempt to pretrain sign language generation models through a cross-modal and multi-level strategy that jointly predicts sign and text tokens, latent features, and 3D sign motions, thereby effectively injecting sign-aware knowledge into MDLMs.
@@ 3 Methodology
An overview of MaDiS is illustrated in Fig.˜2. In the first stage, both text and sign tokens are randomly masked according to a masking ratio $t$, and the MDLM is pretrained with objectives defined across multiple representation spaces (token, latent, and physical) to exploit the benefits from multi-level sign representations. In the subsequent fine-tuning stage, the MDLM keeps text tokens unmasked and learns to generate sign tokens using the proposed unmasking with temporal checkpoints (UTC) strategy and the mixture-of-parts (MoP) embedding layer.
@@ 3.1 Preliminaries
Sign Tokenizer. We adopt the open-sourced SOTA decoupled tokenizer [zuo2025soke] to discretize sign motions, which can be reconstructed with high accuracy. It consists of three parallel VQ-VAEs [vqvae] responsible for different body parts. Each VQ-VAE includes an encoder $\mathcal{E}_{p}$, a decoder $\mathcal{D}_{p}$, and a codebook $\mathcal{C}_{p}=\{{\mathbf{c}_{p}^{i}}\}_{i=1}^{N_{c}}$, where $p\in\{b,l,r\}$ represents the upper body, left hand, and right hand, respectively, and $N_{c}$ is the codebook size. Using this tokenizer, a $T$-frame sign motion sequence $\mathbf{S}\in\mathbb{R}^{T\times d_{s}}$, where $d_{s}=133$ denotes the number of SMPL-X parameters [smplx], is decomposed as three part-wise token sequences $x_{p}=\{x_{p}^{i}\}_{i=1}^{L_{s}}$, where $x_{p}^{i}\in\{1,\dots,N_{c}\}$ and $L_{s}$ denotes the token sequence length.
Masked Diffusion Language Model. MDLMs [llada, smdm] introduce a new language modeling paradigm based on forward-reverse discrete diffusion. In the forward process, the original token sequence $x_{0}$ is corrupted with mask tokens $|\mathrm{M}|$ according to the time (noise level) $t\in[0,1]$. The conditional distribution of the masked sequence $x_{t}$ is defined as $q_{t|0}(x_{t}|x_{0})=\prod_{i=1}^{L}q_{t|0}(x_{t}^{i}|x_{0}^{i})$, where
@@
A rigorous proof in [ou2025your] shows that $q_{0|t}(x_{s}^{i}|x_{t})=p_{\text{data}}(x_{0}^{i}|x_{t}^{\mathrm{UM}})$, where $x_{t}^{\mathrm{UM}}$ denotes the unmasked tokens in $x_{t}$ that are identical to those in $x_{0}$. This implies that the time condition $t$ can be omitted from the model inputs, leading to a simplified implementation of MDLM in which the model only needs to predict masked tokens conditioned on the unmasked ones.
@@ 3.2 Tri-Level Cross-Modal Pretraining
Existing works [chang2024large, tirumala2022memorization] reveal that the knowledge of LMs mainly comes from pretraining. As shown in Fig.˜2, we pretrain the MDLM at three spaces, enabling it to absorb knowledge from sign representations of different levels.
Token Space. At the first level, we pretrain the MDLM by predicting masked tokens. Following [llada], at each training iteration, we uniformly sample a noise level (mask ratio) $t\sim U(0,1)$. Denoting text tokens as $x_{e}=\{x_{e}^{i}\}_{i=1}^{L_{e}}$ and an end-of-sentence token as $|\mathrm{eos}|$, the input sequence is formed by concatenating text tokens, $|\mathrm{eos}|$, and sign tokens: $x_{0}=\mathrm{cat}(x_{e},|\mathrm{eos}|,x_{\{p\}},|\mathrm{eos}|)=\{x_{0}^{i}\}_{i=1}^{L_{e}+L_{s}+2}$. For simplicity, we use $\{p\}$ to denote all part-wise sign tokens, since they are masked and unmasked simultaneously once their index is selected. Each index $i$ is selected with a probability of $t$, and its corresponding token is masked to obtain $x_{t}$. The token-space loss is defined as a cross-entropy loss [llada]:
$\mathcal{L}_{tok}=-\mathbb{E}_{t,x_{0},x_{t}}\left[\frac{1}{t}\sum_{i\in I_{m}}\log p_{\theta}(x_{0}^{i}\mid x_{t})\right],\vskip-8.53581pt$ (3)
@@
Latent Space. Inspired by the effectiveness of joint-embedding predictive architectures (JEPAs) [jepa, i-jepa] in numerous vision tasks [v-jepa, v-jepa2, dino-wm], we extend this idea to MDLM pretraining. Unlike the original JEPAs that train an additional target encoder to extract embeddings, we directly leverage the well-trained codebooks $\mathcal{C}_{p}$, whose embeddings can accurately reconstruct sign motions via the decoders, as the prediction targets to simplify model design.
More specifically, let the last hidden states of the MDLM corresponding to masked sign tokens be $\mathbf{H}\in\mathbb{R}^{|I_{s}|\times d}$, where $I_{s}$ denotes the index set of masked sign tokens. We use three MLPs to map $\mathbf{H}$ into the spaces of the codebooks: $\mathbf{H}_{p}=\mathrm{MLP}_{p}(\mathbf{H})\in\mathbb{R}^{|I_{s}|\times d_{c}}$, where $d_{c}$ denotes the dimension of each code embedding. The target embeddings are then collected from the corresponding code embeddings: $\mathbf{\hat{H}}_{p}=\mathrm{cat}(\{\mathbf{c}^{n_{i}}_{p}\mid i\in I_{s},1\leq n_{i}\leq N_{c}\})$, where $n_{i}$ denotes the code index of the $i$-th token. Finally, the MDLM learns to predict the code embeddings with a smoothed L1 loss ($\|\cdot\|_{1}$): $\mathcal{L}_{lat}=\sum_{p}\|\mathbf{H}_{p}-\mathbf{\hat{H}}_{p}\|_{1}$.
Physical Space. Inspired by the visual-language nature of sign languages, we incorporate 3D sign motion reconstruction as an additional pretraining objective in the physical space. More specifically, we reuse the well-trained VAE decoder $\mathcal{D}_{p}$, which is frozen and used to reconstruct sign motions from the mapped hidden states $\mathbf{H}_{p}$. Denoting the ground-truth part-wise sign motions at the masked indices as $\mathbf{\hat{S}}_{p}$, the objective function in the physical space is defined as: $\mathcal{L}_{phy}=\sum_{p}\|\mathcal{D}_{p}(\mathbf{H}_{p})-\mathbf{\hat{S}}_{p}\|_{1}$.
Unlike existing works [hu2023signbert+, zhou2025scaling, rust-etal-2024-privacy] that mask and reconstruct sign poses, we mask tokens but reconstruct poses (motions), which is more challenging and encourages the model to learn more grounded sign representations. The overall pretraining loss is defined as $\mathcal{L}_{pre}=\mathcal{L}_{tok}+\mathcal{L}_{lat}+\mathcal{L}_{phy}$.